# Small Neutrino Masses from a Decoupled Singlet Scalar Field

### v2

@@ Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$. After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops. We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
@@ I Introduction
In this article, we propose a novel model with a $Z_{4}$ discrete symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate. The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
Taking the decoupling limit for making the singlet scalar field $S$ heavy, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
@@ IV Decoupled scalar and small neutrino masses
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km , and the same results can be obtained from the decoupling limit of the general results in eq. (53) in the appendix. Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq. (4) gets simplified to
@@
In the case with $m_{0}\gg M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$. But, there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
@@ VII Conclusions
We presented a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level. We showed that the spurion scalar opens up a new parameter space for the correct relic density for dark matter and a successful leptogenesis is achieved.

### v4

@@ Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$. After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops for relatively light new fields. We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
@@ I Introduction
In this article, we propose a novel model with a $Z_{4}$ discrete symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, only after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, so neutrino masses are generated at one-loops. Moreover, the lightest neutral dark scalar in the model becomes a dark matter candidate. The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
Taking the decoupling limit where the singlet scalar field $S$ becomes heavier than the $Z_{4}$ breaking scale, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. In the present case, we argue that large masses of extra fields or small Yukawa couplings are not necessary, unlike the case with the $Z_{2}$ symmetry only. We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
@@ IV Decoupled scalar and small neutrino masses
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km , and the same results can be obtained from the decoupling limit of the general results in eq. (53) in the appendix. Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gtrsim m_{0}$, the above expression with eq. (4) gets simplified to
@@
In the case with $m_{0}\gtrsim M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$. But, there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
@@ VII Conclusions
We presented a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level.
Although the $Z_{2}$ symmetry is sufficient for stability of dark matter, its extension to the $Z_{4}$ symmetry in our model allows for explaining the origin of small neutrino masses, even if inert doublet scalar and RH neutrinos are relatively light and the Yukawa couplings for neutrinos are sizable. This option gives rise to phenomenologically a novel window for new particle searches. Moreover, we showed that the spurion scalar responsible for the $Z_{4}$ symmetry breaking opens up a new parameter space for obtaining the correct relic density for inert doublet-like dark matter and a successful leptogenesis is achieved.